# Existence, uniqueness, and regularity of

### v1

@@ 3.2 Stability Theorem
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\leq 0\text{ a.e. in }\mathcal{Y}.$
For all $n\in\mathbb{N}$, $u^{n}\geq g^{n}$ on $\bar{\mathcal{Y}}$ by the comparison principle (Proposition 1). To see this, observe that, for all $n\in\mathbb{N}$, $g^{n}$ is a viscosity subsolution of (Eqn) and $u^{n}$ is an $L^{p}$-supersolution of (Eqn) on $\mathcal{Y}^{n}$. Moreover, the operator $F^{n}$ satisfies Assumptions 4 and 8, and, on $\partial\mathcal{Y}^{n}$, $u^{n}=b^{n}\geq g^{n}$. Thus, Proposition 1 applies, and $u^{n}\geq g^{n}$ on $\bar{\mathcal{Y}^{n}}$. On $\bar{\mathcal{Y}}\setminus\bar{\mathcal{Y}^{n}}$, $u^{n}=b^{n}$ and $b^{n}\geq g^{n}$ by assumption. So, $u^{n}\geq g^{n}$ on $\bar{\mathcal{Y}}$. Passing to the limit, $u\geq g$.
It is therefore enough to show that
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. in }{\mathcal{Y}},$
@@
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. in }{\mathcal{O}},$
Let $\mathcal{O}$ be a strict bounded subset of $\mathcal{Y}$. Since $\mathcal{Y}=\bigcup_{n\in\mathbb{N}}\mathcal{Y}^{n}$, compactness implies that there exists some integer $N\geq 1$ that contains $\mathcal{O}$. Moreover, since the sequence $\{\mathcal{Y}^{n}\}_{n\geq 1}$ is increasing, $\mathcal{O}\subset\mathcal{Y}^{n}$ for all $n\geq N$. Without loss of generality, we assume that $N=1$. For all $m\in\mathbb{N}$, define
As $u^{n}_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right)\leq 0$, for $n\geq m$ and for almost every $(t,x)\in\mathcal{O}$,
$\underline{H}^{m}\left(t,x,u^{n}_{t},u^{n}_{xx}\right)\leq 0.$
Observe also that, since the $F^{n}$’s satisfy Assumptions 4 and 5, for all $m\in\mathbb{N}$ and all $v_{t}\in\mathbb{R}$, $\underline{H}^{m}(t,x,v_{t},\cdot)$ is Lipschitz continuous in $v_{xx}$ a.e. in $\mathcal{O}$ with
Moreover, for any $v_{xx}\in\mathbb{S}^{d}$, by the structure condition (SC), $\underline{H}^{m}$ is bounded in $L^{p}\left(\mathcal{O}\right)$ as $\left(u^{n}\right)_{n\in\mathbb{N}}$ is uniformly bounded in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$ (as a weakly convergent sequence). Finally, by assumption, $u^{n}\rightharpoonup u$ in $W^{1,2,p}\left(\mathcal{O}\right)$. So, Theorem 4.2.6 in Krylov, (2018) implies that, for all $m\in\mathbb{N}$,
$\underline{H}^{m}\left(t,x,u_{t},u_{xx}\right)=\underset{n\geq m}{\inf}\left\{u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\right\}=\underset{n\to\infty}{\lim}\,\underline{H}^{m}\left(t,x,u^{n}_{t},u^{n}_{xx}\right)\leq 0\text{ a.e. in }\mathcal{O}.$
Moreover, $u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\to u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)$ pointwise a.e. in $\mathcal{Y}$ by assumption. Letting $m\to\infty$, we obtain
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. in }\mathcal{O}.$
This shows that $u$ is an $L^{p}$-supersolution of (Eq).
Next, we prove that if the ${u^{n}}$’s are $L^{p}$-subsolutions of (Eqn), then $u$ is an $L^{p}$-subsolution of (Eq).
Note that, for all $n\in\mathbb{N}$, $u^{n}\leq b^{n}$ on $\partial\mathcal{Y}$. Taking limits, $u\leq b$ on $\partial\mathcal{Y}$.
Next, we show that
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\geq 0\text{ a.e. in }\mathcal{Y},$
which holds if and only if
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\geq 0\text{ a.e. in }\mathcal{O},$
@@
Let $\mathcal{O}$ be a strict bounded subset of $\mathcal{Y}$. Since $\mathcal{Y}=\bigcup_{n\in\mathbb{N}}\mathcal{Y}^{n}$ and the $\mathcal{Y}^{n}$’s are an increasing sequence of subsets of $\mathcal{Y}$, there exists $N\in\mathbb{N}$ such that $\mathcal{O}\subset\mathcal{Y}^{n}$, for all $n\geq N$. Without loss of generality, assume that $N=1$. It is then enough to show that, on =$\mathcal{E}=\left\{(t,x)\in\mathcal{O}\,:\,u(t,x)>g(t,x)\right\}$, we have
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\geq 0\text{ a.e. in }\mathcal{O}.$
In particular, if $\mathcal{E}=\emptyset$, we are done. So, suppose not, and let $k$ be large enough so
@@
Then, there exists $N^{k}\in\mathbb{N}$ such that $g^{n}<u^{n}$ on $\mathcal{E}^{k}$ for all $n\geq N^{k}$. To see this, choose $N^{k}$ such that $\left|u^{n}-u\right|<\frac{1}{3k}$ on $\mathcal{E}^{k}$ (which exists since $u^{n}\to u$ uniformly). Then recall that either (i) $g^{n}\uparrow g$ pointwise, or (ii) $g^{n}\to g$ uniformly. So, either (i) $g^{n}\leq g<u-\frac{1}{k}<u-\frac{1}{3k}<u^{n}$ on $\mathcal{E}^{k}$ and we are done, or (ii) $g^{n}$ converge uniformly to $g$. In that case, choose $\tilde{N}^{k}$ such that $\left|g^{n}-g\right|<\frac{1}{3k}$ for all $(t,x)\in\mathcal{O}$, and again, it follows that $g^{n}<g+\frac{1}{3k}<u-\frac{1}{3k}<u^{n}$ in $\mathcal{E}^{k}$ for all $n\geq\max\left\{{N}^{k},\tilde{N}^{k}\right\}$.
@@
$u^{n}_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right)=0\text{ a.e. in }\mathcal{E}^{k}.$
Similarly to the first part of the proof, for all $m\geq N^{k}$, define
@@
Again, observe that, for $n\geq m$ and for almost every $(t,x)\in\mathcal{E}^{k}$,
@@
Moreover, as above, for all $m\in\mathbb{N}$ and all $v_{t}\in\mathbb{R}$, $\bar{H}^{m}(t,x,v_{t},\cdot)$ is Lipschitz continuous in $v_{xx}$ a.e. in $\mathcal{E}^{k}$ with
@@
Also, for any $v_{xx}\in\mathbb{S}^{d}$, by the structure condition (SC), $\bar{H}^{m}$ is bounded in $L^{p}\left(\mathcal{O}\right)$ as $\left(u^{n}\right)_{n\in\mathbb{N}}$ is uniformly bounded in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$ (as a weakly convergent sequence). Finally, by assumption, $u^{n}\rightharpoonup u$ in $W^{1,2,p}\left(\mathcal{E}^{k}\right)$. So, Theorem 4.2.6 in Krylov, (2018) implies, for all $m\in\mathbb{N}$,
$\bar{H}^{m}\left(t,x,u_{t},u_{xx}\right)=\underset{n\geq m}{\inf}\left\{u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\right\}=\underset{n\to\infty}{\lim}\,\bar{H}^{m}\left(t,x,u^{n}_{t},u^{n}_{xx}\right)\geq 0\text{a.e. in }\mathcal{E}^{k}.$
But $u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\to u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)$ pointwise a.e. in $\mathcal{Y}$ by assumption. Letting $m\to\infty$, we obtain
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\geq 0\text{ a.e. in }\mathcal{E}^{k}.$
@@
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\geq 0\text{ a.e. in }\mathcal{E}.$
This concludes the second part of the proof: $u$ is an $L^{p}$-subsolution of (Eq). $\blacksquare$
@@
Theorem 2 is also valid for non-obstacle problems and viscosity solutions. To see this, one can invoke Theorem 6.1 in Crandall et al., (2000) instead of Theorem 4.2.6 in Krylov, (2018) in the proof.
@@
${b}:\bar{\mathcal{Y}}\to\mathbb{R}$ is in $W^{1,2,p}\left(\mathcal{Y}\right)$.
@@
Let $p\in(d+2,\infty]$ and suppose that Assumptions 1’, 2’, and 37 hold. If $\left|A\right|<\infty$, then (1) has an $L^{p}$-solution $u$.
@@
The proof builds on the approximation argument proposed by Byun et al., (2018) and Byun et al., (2022), extending it to fully nonlinear parabolic obstacle problems. We first establish existence and obtain a $W^{1,2,p}$-estimate when the obstacle is in $W^{1,2,p}$ using known existence results and estimates on the regularity of solutions of the Dirichlet problem, e.g., Theorem 12.1.7, 15.1.3, and 15.1.4 in Krylov, (2018). These estimates allow for an arbitrary continuous dependence of the operator on the value of the solution $u$. As a result, we can sidestep the fixed-point argument in Byun et al., (2018), streamlining the proof. We then show by induction that the existence and regularity results for a single $W^{1,2,p}$-obstacle carry to the pointwise maximum of a finite number of $W^{1,2,p}$-obstacles.
@@
We prove Lemma 2 for $p\in(d+2,\infty)$. The case $p=\infty$ follows from the existence of an $L^{p}$-solution for some $p<\infty$ and a classic embedding theorem (see, e.g., Lemma 3.3, page 80, in Ladyženskaja et al., (1968)), which guarantees that $\max\left\{\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\right)},\left\|u_{x}\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\right\}\leq C\left\|u\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}$ for some $C\in\mathbb{R}_{+}$ that depends only on $d,p,diam(\mathcal{X}),T$, and $\left\|\partial\mathcal{X}\right\|_{C^{0,Lip}}$. The structure condition (SC) and equation (8) below then implies that, for almost every $(t,x)\in\mathcal{Y}$,
$\left|u_{t}\right|+\left|u_{xx}\right|\leq C\left(1+\left\|G\right\|_{L^{\infty}\left(\mathcal{Y}\right)}+\underset{a\in A}{\max}\left\|g^{a}\right\|_{W^{1,2,\infty}\left(\mathcal{Y}\right)}+\left\|u\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}\right)$
for some $C=C\left(d,\lambda,\Lambda,R,diam(\mathcal{X}),T,\left\|\partial\mathcal{X}\right\|_{\mathcal{C}^{0,Lip}}\right)\in\mathbb{R}_{+}$. The conclusion follows from estimate (6) and the boundedness of the domain.
@@
Accordingly, consider the case $p\in(d+2,\infty)$. Without loss of generality, assume that $A=\left\{1,\dots,I\right\}$, with $I\in\mathbb{N}$. The proof is by induction on $I$. For all $I\in\mathbb{N}$, define property $P(I)$ as follows:
@@
If $g=\underset{a\in\left\{1,\dots,I\right\}}{\max}\,g^{a}$ with $g^{a}\in W^{1,2,p}\left(\mathcal{Y}\right)$ for all $a\in\left\{1,\dots,I\right\}$, then (1) has a unique $L^{p}$-solution $u\in W^{1,2,p}\left(\mathcal{Y}\right)$. Moreover, $u$ satisfies the estimate (6).
@@
We start by proving our base case: $P(I=1)$. Let $g^{1}\in W^{1,2,p}\left(\mathcal{Y}\right)$ with $g^{1}\leq b$ on $\partial\mathcal{Y}$. For all $\epsilon>0$, let $\Phi_{\epsilon}\in\mathcal{C}^{\infty}\left(\mathbb{R}\right)$ be a nondecreasing function such that $\Phi_{\epsilon}(a)=0$ if $a\leq 0$, and $\Phi_{\epsilon}(a)=1$ if $a\geq\epsilon$. In particular, for all $a\in\mathbb{R}$, $\Phi_{\epsilon}(a)\in[0,1]$. Define also $h^{1}(t,x)=-g^{1}_{t}-F\left(t,x,g^{1},g^{1}_{x},g^{1}_{xx}\right)$. Let $\left(\epsilon_{n}\right)_{n\in\mathcal{N}}\subseteq\mathbb{R}_{++}$ be a sequence such that $\epsilon_{n}\to 0$ as $n\to\infty$; and consider the following auxiliary nonlinear Dirichlet problems. For all $n\in\mathbb{N}$,
@@
$\begin{cases}{F}\left(t,x,u^{n},u^{n}_{t},u^{n}_{x},u^{n}_{xx}\right)=h^{1}(t,x)^{+}\Phi_{\epsilon_{n}}\left(u^{n}-g^{1}\right)-h^{1}(t,x)^{+}\text{ in }\mathcal{Y},\\ u^{n}=b\text{ on }\partial\mathcal{Y}.\end{cases}$ (D)
@@
By Assumptions 4 and 7, $h^{1}(t,x)$ is in ${L}^{p}\left(\mathcal{Y}\right)$ with
$\left\|h^{1}\right\|_{L^{p}(\mathcal{Y})}\leq C^{h^{1}}\left(1+\left\|g^{1}\right\|_{W^{1,2,p}(\mathcal{Y})}+\left\|G\right\|_{L^{p}(\mathcal{Y})}\right),$ (7)
@@
where $C^{h^{1}}=C\left(d,p,\lambda,\Lambda,{R}\right)\in\mathbb{R}_{+}$. By Theorem 15.1.4 in Krylov, (2018), for each $n\in\mathbb{N}$, there exists a solution $u^{n}\in W^{1,2,p}\left(\mathcal{Y}\right)$ of (D). Moreover, by Theorem 15.1.3 (using Lemma 12.1.9 in Krylov, (2018) to bound $\left\|u\right\|_{L^{\infty}\left(\mathcal{Y}\right)}$ and Theorem 9.8.1 in Krylov, (2018) to obtain the modulus of continuity of $u$),
$\left\|u^{n}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}\leq C\left(1+\left\|g^{1}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|b\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}\right),$
@@
where $C=C\left(d,p,\lambda,\Lambda,R,diam(\mathcal{X}),T,\left\|\partial\mathcal{X}\right\|_{\mathcal{C}^{1,Lip}}\right)\in\mathbb{R}_{+}$ is independent of $\epsilon_{n}$.
@@
$W^{1,2,p}(\mathcal{Y})$ is separable and reflexive. [^12] The space ${L}^{p}$ are uniformly convex for $p\in(1,\infty)$, hence so is $W^{1,2,p}\left(\mathcal{Y}\right)$ for $p\in(1,\infty)$. Therefore it is reflexive by Theorem 1.40 in Demengel et al., (2012). Therefore, its closed bounded subsets are weakly sequentially compact by Theorem 1.32 in Demengel et al., (2012). Moreover $W^{1,2,p}\left(\mathcal{Y}\right)$ is compactly embedded in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ by Rellich-Kondrachov’s theorem (Theorem 2.84 in Demengel et al., (2012)). So, there exists a function $u\in W^{1,2,p}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ and a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u^{n}\right)_{n\in\mathbb{N}}$ such that
@@
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. in }\mathcal{Y}.$
@@
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)=0\text{ in }\mathcal{Y}.$
@@
We will show that, for all $j\in\mathbb{N}$, $\mathcal{V}_{n_{j}}$ is empty, by contradiction. Suppose not, i.e., $\mathcal{V}_{n_{j}}\neq\emptyset$ for some $j\in\mathbb{N}$. Since both $u^{n_{j}}$ and $g^{1}$ are continuous on $\mathcal{Y}$, $\mathcal{V}_{n_{j}}$ is open. Moreover, since $b\geq g^{1}$, $u^{n_{j}}=g^{1}$ on $\partial\mathcal{V}_{n_{j}}$. On $\mathcal{V}_{n_{j}}$, $\Phi_{\epsilon_{n_{j}}}\left(u^{{n_{j}}}(t,x)-g(t,x)\right)=0$. So $u^{{n_{j}}}\in W^{1,2,p}\left(\mathcal{Y}\right)$ is an $L^{p}$-solution of
$\begin{cases}u^{n_{j}}_{t}+F\left(t,x,u^{{n_{j}}},u_{x}^{{n_{j}}},u_{xx}^{{n_{j}}}\right)=-{h^{1}}(t,x)^{+}\text{ in }\mathcal{V}_{n_{j}},\\ u^{{n}}=g^{1}\text{ on }\partial\mathcal{V}_{n_{j}}.\end{cases}$

### v2

@@ 4.2 Stability Theorem
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\geq 0\text{ a.e. on }\mathcal{Y},$
@@
$\max\left\{u_{t}+F\left(t,x,u,u_{x},u_{xx}\right),\,g-u\right\}\geq 0\text{ a.e. on }\mathcal{O},$
@@
Let $\mathcal{O}$ be a strict bounded subset of $\mathcal{Y}$. Since $\mathcal{Y}=\bigcup_{n\in\mathbb{N}}\mathcal{Y}^{n}$ and the sequence of subsets $\mathcal{Y}^{n}$ is increasing, there exists $N\in\mathbb{N}$ such that $\mathcal{O}\subset\mathcal{Y}^{n}$, for all $n\geq N$. Without loss of generality, assume that $N=1$. It is then enough to show that, on =$\mathcal{E}=\left\{(t,x)\in\mathcal{O}\,:\,u(t,x)>g(t,x)\right\}$, we have
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\geq 0\text{ a.e. on }\mathcal{O}.$
If $\mathcal{E}=\emptyset$, there is nothing left to prove. Suppose instead that $\mathcal{E}$ is nonempty, and let $k$ be large enough so that
@@
Then, there exists $N^{k}\in\mathbb{N}$ such that $g^{n}<u^{n}$ on $\mathcal{E}^{k}$ for all $n\geq N^{k}$. To see this, choose $N^{k}$ such that $\left|u^{n}-u\right|<\frac{1}{3k}$ on $\mathcal{E}^{k}$ (which exists since $u^{n}\to u$ uniformly). Then recall that either (i) $g^{n}\uparrow g$ pointwise, or (ii) $g^{n}\to g$ uniformly. So, either (i) $g^{n}\leq g<u-\frac{1}{k}<u-\frac{1}{3k}<u^{n}$ on $\mathcal{E}^{k}$ and we are done, or (ii) $g^{n}$ converge uniformly to $g$. In that case, choose $\tilde{N}^{k}$ such that $\left|g^{n}-g\right|<\frac{1}{3k}$ for all $(t,x)\in\mathcal{O}$, and again, it follows that $g^{n}<g+\frac{1}{3k}<u-\frac{1}{3k}<u^{n}$ on $\mathcal{E}^{k}$ for all $n\geq\max\left\{{N}^{k},\tilde{N}^{k}\right\}$.
@@
$u^{n}_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u^{n}_{xx}\right)=0\text{ a.e. on }\mathcal{E}^{k}.$
Similarly to the first part of the proof, define for all $m\geq N^{k}$
@@
We note once more that, for $n\geq m$ and for almost every $(t,x)\in\mathcal{E}^{k}$,
@@
Moreover, as above, for all $m\in\mathbb{N}$ and all $v_{t}\in\mathbb{R}$, $\bar{H}^{m}(t,x,v_{t},\cdot)$ is Lipschitz continuous in $v_{xx}$ a.e. on $\mathcal{E}^{k}$ with
@@
Also, for any $v_{xx}\in\mathbb{S}^{d}$, by the structure condition (SC), $\bar{H}^{m}$ is bounded in $L^{p}\left(\mathcal{O}\right)$ because the functions $u^{n}$, $n\in\mathbb{N}$, are uniformly bounded in $W^{1,2,p}_{loc}\left(\mathcal{Y}\right)$ as they form a weakly convergent sequence. Finally, by assumption, $u^{n}\rightharpoonup u$ in $W^{1,2,p}\left(\mathcal{E}^{k}\right)$. Theorem 4.2.6 in Krylov, (2018) then implies for all $m\in\mathbb{N}$ that
$\bar{H}^{m}\left(t,x,u_{t},u_{xx}\right)=\underset{n\geq m}{\inf}\left\{u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\right\}=\underset{n\to\infty}{\lim}\,\bar{H}^{m}\left(t,x,u^{n}_{t},u^{n}_{xx}\right)\geq 0\text{a.e. on }\mathcal{E}^{k}.$
By assumption, $u_{t}+F^{n}\left(t,x,u^{n},u^{n}_{x},u_{xx}\right)\to u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)$ pointwise a.e. on $\mathcal{Y}$. Letting $m\to\infty$, we obtain
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\geq 0\text{ a.e. on }\mathcal{E}^{k}.$
@@
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\geq 0\text{ a.e. on }\mathcal{E}.$
This shows that $u$ is an $L^{p}$-subsolution of (Eq). $\blacksquare$
@@
Theorem 2 is also valid for non-obstacle problems, as can be seen from the proof. Moreover, it also holds for viscosity solutions. To see this, one can invoke Theorem 6.1 in Crandall et al., (2000) instead of Theorem 4.2.6 in Krylov, (2018) in the proof.
@@
${b}:\bar{\mathcal{Y}}\to\mathbb{R}$ is in $W^{1,2,p}\left(\bar{\mathcal{Y}}\right)$.
@@
Suppose that Assumptions 1’, 2’, and 37 hold for some $p\in(d+2,\infty)$. [^17] The parameter $p$ appears in Assumptions 3 and 7. If $\left|A\right|<\infty$, then (1) has an $L^{p}$-solution $u$.
@@
The proof builds on the approximation argument proposed by Byun et al., (2018) and Byun et al., (2022), extending it to fully nonlinear parabolic obstacle problems. Byun et al., (2018) approximate the obstacle problem by a Dirichlet problem similar to (Dn). They then use the Schauder fixed-point theorem to show the existence of $L^{p}$-solutions whose norms are bounded by a constant independent of $\epsilon$. Instead, we appeal to known existence results and estimates on the regularity of solutions of the Dirichlet problem, e.g., Theorem 12.1.7, 15.1.3, and 15.1.4 in Krylov, (2018) that allow for an arbitrary continuous dependence of the operator on the value of the solution $u$. As a result, we can sidestep the fixed-point argument in Byun et al., (2018), streamlining the proof of the existence of a solution and $W^{1,2,p}$-estimate when the obstacle is in $W^{1,2,p}$. We then show by induction that the existence and regularity results for a single $W^{1,2,p}$-obstacle carry to the pointwise maximum of a finite number of $W^{1,2,p}$-obstacles.
@@
Without loss of generality, assume that $A=\left\{1,\dots,I\right\}$, for some integer $I\geq 1$. The proof proceeds by induction on $I$. For all $I\geq 1$, define property $\bf{P(I)}$ as follows:
@@
If $g=\underset{a\in\left\{1,\dots,I\right\}}{\max}\,g^{a}$ with $g^{a}\in W^{1,2,p}\left(\mathcal{Y}\right)$ for all $a\in\left\{1,\dots,I\right\}$, then (1) has a $L^{p}$-solution $u\in W^{1,2,p}\left(\mathcal{Y}\right)$. Moreover, $u$ satisfies the estimate (6).
@@
We start by proving the base case $\bf{P(1)}$. Thus, let $g^{1}\in W^{1,2,p}\left(\mathcal{Y}\right)$ be such that $g^{1}\leq b$ on $\partial\mathcal{Y}$. Let $h^{1}(t,x)=-g^{1}_{t}-F\left(t,x,g^{1},g^{1}_{x},g^{1}_{xx}\right)$. For each $\epsilon>0$, let $\Phi_{\epsilon}\in\mathcal{C}^{\infty}\left(\mathbb{R}\right)$ be a nondecreasing function such that $\Phi_{\epsilon}(a)=0$ if $a\leq 0$, and $\Phi_{\epsilon}(a)=1$ if $a\geq\epsilon$. In particular, for all $a\in\mathbb{R}$, $\Phi_{\epsilon}(a)\in[0,1]$. Finally, consider a sequence $\left(\epsilon_{n}\right)_{n\in\mathcal{N}}\subseteq\mathbb{R}_{++}$ such that $\epsilon_{n}\to 0$ as $n\to\infty$, and consider the following auxiliary nonlinear Dirichlet problem for each $n\in\mathbb{N}$:
@@
$\begin{cases}{F}\left(t,x,u^{n},u^{n}_{t},u^{n}_{x},u^{n}_{xx}\right)=h^{1}(t,x)^{+}\Phi_{\epsilon_{n}}\left(u^{n}-g^{1}\right)-h^{1}(t,x)^{+}\text{ on }\mathcal{Y},\\ u^{n}=b\text{ on }\partial\mathcal{Y}.\end{cases}$ (Dn)
@@
Here $h^{1}(t,x)^{+}=\max\left\{h^{1}(t,x),0\right\}$ stands for the nonnegative part of $h^{1}$. By Assumptions 4 and 7, $h^{1}(t,x)$ is in ${L}^{p}\left(\mathcal{Y}\right)$ and
$\left\|h^{1}\right\|_{L^{p}(\mathcal{Y})}\leq C^{h^{1}}\left(1+\left\|g^{1}\right\|_{W^{1,2,p}(\mathcal{Y})}+\left\|G\right\|_{L^{p}(\mathcal{Y})}\right)$ (7)
@@
where $C^{h^{1}}=C\left(d,p,\lambda,\Lambda,{R}\right)\in\mathbb{R}_{+}$. By Theorem 15.1.4 in Krylov, (2018), for each $n\in\mathbb{N}$, there exists a solution $u^{n}\in W^{1,2,p}\left(\mathcal{Y}\right)$ of (Dn). Moreover, by Theorem 15.1.3 in the same work,
$\left\|u^{n}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}\leq C\left(1+\left\|h^{1}\right\|_{L^{p}(\mathcal{Y})}+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}+\left\|b\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|u^{n}\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\right)$
@@
where $C=C\left(d,p,\lambda,\Lambda,R,diam(\mathcal{X}),T,\left\|\partial\mathcal{X}\right\|_{\mathcal{C}^{1,Lip}}\right)\in\mathbb{R}_{+}$. By Lemma 12.1.9 in Krylov, (2018),
$\left\|u^{n}\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\leq C\left(1+\left\|h^{1}\right\|_{L^{p}(\mathcal{Y})}+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}+\left\|b\right\|_{L^{\infty}\left(\partial\mathcal{Y}\right)}\right),$
where $C=C\left(d,p,\lambda,\Lambda,R,diam(\mathcal{X}),T\right)\in\mathbb{R}_{+}$. Finally, $\left\|b\right\|_{L^{\infty}\left(\partial\mathcal{Y}\right)}\leq$$\left\|b\right\|_{L^{\infty}\left(\mathcal{Y}\right)}\leq C\left\|b\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}$, where $C=C(d,p,diam(\mathcal{X}),T,\left\|\partial\mathcal{X}\right\|_{\mathcal{C}^{0,Lip}})$, by the Morrey-Sobolev embedding theorem. Combining the inequalities above, we obtain
$\left\|u^{n}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}\leq C\left(1+\left\|g^{1}\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|b\right\|_{W^{1,2,p}\left(\mathcal{Y}\right)}+\left\|G\right\|_{L^{p}\left(\mathcal{Y}\right)}\right)$
where $C=C\left(d,p,\lambda,\Lambda,R,diam(\mathcal{X}),T,\left\|\partial\mathcal{X}\right\|_{\mathcal{C}^{1,Lip}}\right)\in\mathbb{R}_{+}$ is independent of $n$.
@@
$W^{1,2,p}(\mathcal{Y})$ is separable and reflexive. [^18] The space ${L}^{p}$ is uniformly convex for $p\in(1,\infty)$, hence so is $W^{1,2,p}\left(\mathcal{Y}\right)$. Therefore, this latter space is reflexive by Theorem 1.40 in Demengel et al., (2012). Therefore, its closed bounded subsets are weakly sequentially compact by Theorem 1.32 in Demengel et al., (2012). Moreover $W^{1,2,p}\left(\mathcal{Y}\right)$ is compactly embedded in $\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ by the Rellich-Kondrachov theorem (see Theorem 2.84 in Demengel et al., (2012)). Therefore, there exists a function $u\in W^{1,2,p}\left(\mathcal{Y}\right)\cap\mathcal{C}^{0}\left(\bar{\mathcal{Y}}\right)$ and a subsequence $\left(u^{n_{j}}\right)_{j\in\mathbb{N}}\subseteq\left(u^{n}\right)_{n\in\mathbb{N}}$ such that
@@
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)\leq 0\text{ a.e. on }\mathcal{Y}.$
@@
$u_{t}+F\left(t,x,u,u_{x},u_{xx}\right)=0\text{ on }\mathcal{Y}.$
@@
We will show that $\mathcal{V}_{n_{j}}$ is empty for each $j\in\mathbb{N}$. Suppose by way of contradiction that $\mathcal{V}_{n_{j}}\neq\emptyset$ for some $j\in\mathbb{N}$. Since both $u^{n_{j}}$ and $g^{1}$ are continuous on $\mathcal{Y}$, $\mathcal{V}_{n_{j}}$ is open. Moreover, since $b\geq g^{1}$, $u^{n_{j}}=g^{1}$ on $\partial\mathcal{V}_{n_{j}}$. On $\mathcal{V}_{n_{j}}$, $\Phi_{\epsilon_{n_{j}}}\left(u^{{n_{j}}}(t,x)-g(t,x)\right)=0$. So $u^{{n_{j}}}\in W^{1,2,p}\left(\mathcal{Y}\right)$ is an $L^{p}$-solution of
$\begin{cases}u^{n_{j}}_{t}+F\left(t,x,u^{{n_{j}}},u_{x}^{{n_{j}}},u_{xx}^{{n_{j}}}\right)=-{h^{1}}(t,x)^{+}\text{ on }\mathcal{V}_{n_{j}},\\ u^{n_{j}}=g^{1}\text{ on }\partial\mathcal{V}_{n_{j}}.\end{cases}$